Michael gets paid £y per week for his paper round after four weeks he gets paid but £2 is deducted for lateness. write an expression for the amount Michael gets. when he opens his pay packet he finds he has been paid £22 write and solve an equation to find out how much Michael is paid per week
Answer by math_tutor2020(3816)


y = amount he is paid per week (in pounds)
4y = amount he is paid over 4 weeks total, ignoring the penalty fee
4y-2 = now accounting for the 2 pound deduction

The amount Michael is paid is 4y-2 pounds.
He finds out he's actually paid 22 pounds.
Set those two expressions equal to one another and solve.

4y-2 = 22
4y = 22+2
4y = 24
y = 24/4
y = 6
Michael is paid 6 pounds per week

If he wasn't penalized, then he'd get paid 4y = 4*6 = 24 pounds over the four week period.
However, he's penalized those 2 pounds to get paid 24-2 = 22 pounds instead.
This matches with what your teacher gave you, so it confirms the answer.